# Reasoning in Token Economies: Budget-Aware Evaluation of LLM Reasoning Strategies

### v2

@@ 1 Introduction
Through this research, we offer a more nuanced understanding of reasoning strategies in LLMs by considering both their performance and computational costs.
@@
We explore the dynamics of reasoning strategies, highlighting that MAD underperforms as diversity diminishes with each additional round. Conversely, Self-Consistency excels due to the independence of samples boosting diversity and its effectiveness in scenarios where the likelihood of a correct answer exceeds 50%.
We conduct ablation study on ToT and Reflexion by segregating budget into answer generation and evaluator budget. We found that self-evaluation has impact but is model and dataset-dependent. We verify this finding by explicitly testing the self-evaluation ability of GPT-4 on five datasets as well as utilizing our newly proposed strategy .
@@ 5.3 Reflexion ablation study
5.4 Self-evaluation effectiveness is dataset dependent
In this section, we investigate further how a self-evaluator can benefit the reasoning process in a budget-aware setting. First, we benchmark self-evaluation ability and calibration results on the MATH dataset. Then we do independent studies to demonstrate that self-evaluation is correlated to dataset, confirming results from the last section. Finally, we incorporate self-evaluation to propose a new reasoning strategy that improves upon reasoning ability on two math reasoning datasets.
5.4.1 Accuracy and Calibration of Self-Evaluation
@@
5.4.2 Self-evaluation ability on different datasets
Table 1 shows the self-evaluation accuracy for GPT-4 for multiple datasets. The self-evaluation accuracy turns out to be heavily dependent on the dataset. It is possible that for problems that are too hard for the model, it ends up weighing the writing style of the answer much more heavily than the correctness of all the intermediate steps when doing the evaluation. We saw empirical evidence in Figure 15 where easier datasets are better calibrated and vice versa. We examine this in more detail in the appendix D, which will also include how we specifically do self-evaluation and how well-calibrated they are.
5.4.3 Self-Confident Self-Consistency ()
Although the self-evaluation incurs additional costs to the reasoning strategy, plenty of evidence including some past work demonstrates the usefulness of self-evaluation (Olausson et al., 2023; Li et al., 2023). We thus propose to leverage that ability by weighting the SC by the confidence the model has in its answer, derived via the Binary evaluation strategy. We call this score the Self-Confident Self-Consistency () score. We showed that beats self-consistency on 2 out of 3 math reasoning tasks and is at least as good as self-consistency on others with very similar budgets shown in Figure 11. We can observe that datasets that GPT-4 has a decent self-evaluation performance on, would also have a decent performance. This further proves that self-evaluation is dataset dependent. More details about the implementation and results can be found at Appendix D.3.
@@ 6 Conclusion
In this paper, we examined the performance of seven reasoning strategies with respect to the often overlooked metric of budget. We used budget metrics of number of queries and number of tokens to reflect the different ways LLMs are used (querying LLM APIs and self-hosting models). We identified self-evaluation as an important aspect of multiple reasoning strategies and analyzed different prompting strategies to have the model evaluate its generations. We then evaluated and benchmark self-evaluation and found that self-evaluation performance is model and dataset dependent. We further proved this by introducing a new variant of SC weighted by self-evaluation which showed gains for math reasoning tasks. We also analyzed the answer generation and evaluation budget separately finding that in budget-constrained situations, it can be sometimes beneficial to use a weaker evaluator and still be able to obtain most of the performance one would get with a stronger evaluator. With the current popularity of reasoning strategies, we think this more balanced budget-aware metric is beneficial for the community and helps set the correct trajectory for future LLM research.
@@ D.1 Self-Evaluation Method
Binary222We also investigate a variant where we ask the model to think step by step before evaluating. While we see a small increase in performance for such a strategy, it also necessitates a big increase in the token budget. Further analysis is in the Supplement. - we ask the model to output Yes/No as to whether the answer is correct. We do this multiple times and take the fraction of times the model answers Yes as the confidence of the model in the answer.

### v3

@@ 1 Introduction
We explore the dynamics of reasoning strategies, highlighting that MAD underperforms as diversity diminishes with each round. Conversely, Self-Consistency excels due to the independence of samples boosting diversity and its effectiveness in scenarios where the likelihood of being correct exceeds 50%.
We conduct ablation studies on ToT and Reflexion by segregating the budget into answer generation and evaluation budgets. We found that self-evaluation is promising at increasing performance while being cost-effective but currently LLM can’t self-evaluate well.
@@ 5.3 Reflexion ablation study
5.4 Self-evaluation is a promising budget-efficient improvement but is currently lacking
Self-evaluation usually involves very few tokens generated since evaluation is short. This can be potentially very cost-effective since prefilling is cheaper and faster. In this section, we investigate further how a self-evaluator can benefit the reasoning process in a budget-aware setting and demonstrate why there may still be a long way to go. In this section, we investigate further how a self-evaluator can benefit the reasoning process in a budget-aware setting.
Accuracy and Calibration of Self-Evaluation
@@ 5.4 Self-evaluation is a promising budget-efficient improvement but is currently lacking
Self-evaluation ability on different datasets
Table 1 shows the self-evaluation accuracy for GPT-4 for multiple datasets. The self-evaluation accuracy turns out to be heavily dependent on the dataset. It is possible that for problems that are too hard for the model, it ends up weighing the writing style of the answer much more heavily than the correctness of all the intermediate steps when doing the evaluation. We saw empirical evidence in Figure 15 where easier datasets are better calibrated and vice versa. On harder tasks like TheoremQA, GPT-4’s accuracy is close to random. This means LLMs have a long way to go before they are reliable evaluators. We examine this in more detail in the appendix D.
Self-Confident Self-Consistency ()
As an investigation of using self-evaluation to improve reasoning procedure, we propose to weigh the SC by the confidence the model has in its answer, derived from self-evaluation. We call this score the Self-Confident Self-Consistency () score. We showed that beats self-consistency on GSM8k and MATH while fall behind on the other three as shown in Figure 11.222More details can be found at Appendix D.3. This shows that although theoretically, self-evaluation is promising (shown with oracle results in Figure 10), it is still lacking in practice due to low accuracy.
@@ 6 Conclusion
In this paper, we examined the performance of seven reasoning strategies on the often overlooked metric of budget. We used budget metrics of queries and tokens to reflect various ways LLMs are used (LLM APIs or self-host). We identified self-evaluation as an important aspect of many reasoning strategies and analyzed different prompting strategies to have the model evaluate its generations. We then evaluated self-evaluation and found that although self-evaluation could be promising at improving performance while being cost-effective, current LLMs are mostly incapable of doing that. With the current popularity of reasoning strategies, we think this more balanced budget-aware metric is beneficial for the community and helps set the correct trajectory for future LLM research.
@@ D.1 Self-Evaluation Method
Binary333We also investigate a variant where we ask the model to think step by step before evaluating. While we see a small increase in performance for such a strategy, it also necessitates a big increase in the token budget. Further analysis is in the Supplement. - we ask the model to output Yes/No as to whether the answer is correct. We do this multiple times and take the fraction of times the model answers Yes as the confidence of the model in the answer.